FIRST DETECTION OF GROUND STATE ROTATIONAL LINES OF DOUBLY DEUTERATED AMMONIA

\underline{D.C. LIS}, California Institute of Technology, Downs Laboratory of Physics 320-47, Pasadena, CA 91125; M. GERIN, LERMA, UMR 8112 du CNRS, Observatoire de Paris and Ecole Normale Superiéure, 24 Rue Lhomond, 75231 Paris cedex 05, France; E. ROUEFF, LUTH and UMR 8102 du CNRS, Observatoire de Paris, Section de Meudon, Place J. Janssen, 92195 Meudon, France; C. VASTEL, and T.G. PHILLIPS, California Institute of Technology, Downs Laboratory of Physics 320-47, Pasadena, CA 91125.

We report the first detection of the N_{K_{A}K_{C}} = 1_{11} → 0_{00} and 1_{10} → 0_{00} ground state rotational lines of o-ND_{2}H at 335.5 and 388.7 GHz, obtained in the Lynds 1689N and Barnard 1 molecular clouds using the Caltech Submillimeter Observatory (CSO). The hyperfine splitting is detected in both lines allowing direct determination of the line opacity and the excitation temperature. The 1_{11} → 0_{00} line of p-ND_{2}H at 335.4 GHz is also detected in both sources and the derived ortho-to-para ratio is consistent with the expected LTE value of 2:1. The two ground state o-ND_{2}H transitions have different spontaneous emission rate coefficients and therefore critical densities. With the help of modeling tools, they can thus provide good constraints on the excitation conditions and gas kinematics in dense, cold pre-stellar and protostellar cores. These places are sites of strong molecular depletion and heavy deuteration. Non-accreting molecules, H_{3}^{+} and its isotopologues, are difficult to study, but in several cases it appears that ammonia and its isotopologues are not completely frozen out. On good sites, such as Mauna Kea or Chajnantor, the submillimeter ND_{2}H lines are easier to detect than the 3 mm lines previously used to study the distribution and abundance of this molecule in the interstellar medium. Interferometric observations of these submillimeter ND_{2}H lines with the Submillimeter Array (SMA), and subsequently the Atacama Large Millimeter Array (ALMA), will provide new opportunities to image the inner structure of cold, dense cores, where most molecules are depleted onto dust grains.